OpenStudy (anonymous):

I need help verifying the following trigonometric identity, csc(-x)-1=(cot^2(x))/csc(-x)+1

OpenStudy (ranga):

multiply both sides by { csc(-x) + 1 } and simplify. Make use of the following identities: (a+b) * (a-b) = a^2 - b^2 and 1 + cot^2(x) = csc^2(x)